# Many-Body Localization in the Age of Classical Computing

### v2

@@ VI.2.2 Quantum information measures
$S_{A}=-\mathrm{tr}(\rho_{A}\log_{2}\rho_{A}),$ (38)
where $\rho_{A}=\mathrm{tr}_{B}\ket{\psi}\bra{\psi}$ | is the reduced density matrix of subsystem $A$, obtained by tracing out the degrees of freedom of subsystem $B$. Quantum mutual information (QMI) [445] measures correlation between subsystems $A_{1}$ and $A_{2}$ of the system, and can be calculated as
@@
Entanglement entropy has been employed in the studies of MBL from its early days [697]. Area-law of the entanglement entropy stipulated for eigenstates in the MBL phase [52, 579]. It predicts that the entanglement entropy $S_{A}$ (38) is proportional to the boundary of the subsystem $A$, which, for 1D systems, implies $S_{A}\sim\mathrm{const}$ independently of the size $\ell_{A}$ of the subsystem $A$, provided that $\ell_{A}\gg 1$. This is consistent with the LIOMs picture, as discussed in Sec. III.4. In contrast, the mid-spectrum eigenstates in ergodic regime are nearly maximally entangled [458, 74], leading to volume-law $S_{A}\sim\ell_{A}$. Global conservation laws, such as $U(1)$ symmetry of (7) affect the constant term in the volume-law scaling of $S_{A}$ [680].
@@
For a meaningful comparison of entanglement entropy for different system sizes, it is convenient to rescale the entanglement entropy by the average $S_{\mathrm{RMT}}$ for Haar random states in the considered subspace of the Hilbert space, and consider $s=S_{A}/S_{\mathrm{RMT}}$. Averaging this quantity over mid-spectrum eigenstates |$\ket{m}$ and disorder realizations, leads to $\overline{s}$ which may serve as an ergodicity breaking indicator. The rescaled entanglement entropy, $\overline{s}$, across the ETH-MBL crossover in the $J_{1}$-$J_{2}$ model converges, with increasing the system size $L$, to unity in the ergodic regime and it decreases towards 00 in the MBL regime, as shown in Fig. 7(c). Even deep in the ergodic regime, we observe a small correction of the order of $\mathcal{O}(1)$ to the value of the averaged entanglement entropy. The magnitude of the correction depends on the choice of the parameters of the model [328]. Notably, we observe a significant drift of the crossing point of $\overline{s}(W)$ curves with the increase of system sizes. This drift is analogous to the behavior of the averaged gap ratio $\overline{r}$ and the fractal dimension $D_{q}$ discussed in the preceding sections, disallowing a single parameter scaling and preventing one from reaching unambiguous conclusions about the fate of the crossover in the $L\to\infty$ limit.
@@
Interestingly, the extent of the drift of the crossing point may depend on the considered observable. For instance, within the range of system sizes considered by us, the drifts of the crossings points are larger for the average gap ratio $\overline{r}$ and the rescaled entanglement entropy $\overline{s}$ than for the fractal dimension $D_{q}$, c.f. Fig. 4, Fig. 6, and Fig. 7, respectively. Similarly, the QMI calculated for extensive subsystems $A_{1}$ and $A_{2}$ of the Kicked Ising model in [596] exhibits a weaker drift of the crossing point than the entanglement entropy. A consistency of the extrapolations to the large $L$ limit of crossing points $W^{*}(L)$ yields consistent results for various observables [597, 594, 596]. The consistency between extrapolations of the crossing points for various observables may suggest their relevance. Nevertheless, such extrapolations remain uncontrolled in absence of solid arguments for the form of the extrapolating functions.
@@ VI.3 Matrix elements of local observables
Matrix elements $\langle n|\hat{A}|m\rangle$ of local observables $\hat{A}$ in eigenstates |$\ket{m}$ of many-body Hamiltonian $\hat{H}$ provide a direct link between the features of system’s spectrum and measurable dynamical properties. Analysis of properties of matrix elements of local observables is an important way of quantifying ergodicity and its breakdown in many-body systems.
@@ VI.3.1 Spectral function
The ETH ansatz for $\langle n|\hat{A}|m\rangle$ (3) consists of the diagonal term ($m=n$) and the off-diagonal terms ($m\neq n$). We employ a spectral function of a local observable $\hat{A}$ as an organizing principle for studies of both diagonal and off-diagonal matrix elements of $\hat{A}$. The spectral function is defined as the Fourier transform of the autocorrelation function $\left\langle A(t)A\right\rangle$, i.e.,
@@
where the average $\langle$.$\rangle$ . ⟩ is taken with respect to a selected ensemble of states. In the following, to stay in a direct correspondence with the quantities analyzed so far in this section, we focus on microcanonical ensemble of states $\mathcal{M}_{E_{c}}$ around the middle of the spectrum. In that case, calculating the Fourier transform in (40), we arrive at
@@
where $\mathcal{M}_{E_{c}}=\{m:|E_{m}-E_{c}|<\delta E\}$ defines the microcanonical energy shell around the target energy $E_{c}$ (taken to be the arithmetic average of the ground state and highest excited state energy), and $\delta E$ decreases with $L$, so that the number of states $|\mathcal{M}_{E_{c}}|$ in the energy shell is no bigger than a fixed number $n_{\mathrm{ev}}$ of states (we take $n_{\mathrm{ev}}=\min\{400,\mathcal{N}/20\}$), where $\mathcal{N}$ is the Hilbert space dimension. 212121 We note that various definitions of $f^{2}(\omega)$ can be found in the literature. The average $\langle$.$\rangle$ . ⟩ may be taken with respect to the infinite temperature ensemble of states in [572, 678]. The spectral function considered by [423, 299, 560] involves contributions of eigenstates |$\ket{m},\ket{n}$ with the average energy fixed at a selected target value, while [70] assumes that both eigenstates |$\ket{m},\ket{n}$ belong to a selected energy interval.
@@
The $f^{2}(\omega)$ function is directly linked with the dynamics of the system as the Fourier transform of the autocorrelation function. The spectral function enters the fluctuation-dissipation relations [310], and, if the operator $\hat{A}$ is a density of a conserved quantity, e.g., a spin current operator [502], $f^{2}(\omega)$ is related to the conductivity of the system [8]. Comparison of (41) with the ETH ansatz provides a link between the spectral function and the structure function $f_{\mathcal{A}}(\bar{E},\omega_{mn})$ in (3). Moreover, the small frequency behavior of the spectral function, $f^{2}(\omega)\stackrel{{\scriptstyle\omega\ll 1}}{{=}}C_{0}\delta(\omega)$, reflects the averaged square of the diagonal matrix element of $\hat{A}$, $C_{0}=\frac{1}{|\mathcal{M}_{E_{c}}|}\sum_{m\in\mathcal{M}_{E_{c}}}|\bra{m}\hat{A}\ket{m}|^{2}$, since the length of the energy interval corresponding to $\mathcal{M}_{E_{c}}$ vanishes in the large $L$ limit.
@@
Figure 8: The spectral function $f^{2}(\omega)$ calculated for $\hat{A}=\hat{S}^{z}_{j}$ in the $J_{1}$-$J_{2}$ spin chain of size $L$ in the ergodic regime, (a), for $W=1.5$ and in the MBL regime, (b), for $W=12$. The Heisenberg energy $\omega_{\mathrm{H}}\sim e^{-cL}$ is denoted by red squares, while the Thouless energy $\omega_{\mathrm{Th}}$ by red circles. The dashed line denotes a diffusive tail $f^{2}(\omega)\propto\omega^{-1/2}$. The stiffness $C_{0}$ as function of disorder strength $W$ for various $L$ distinguishes ETH and MBL regimes.
@@
The spectral function $f^{2}(\omega)$ evaluated for the operator $\hat{A}=S^{z}_{j}$, where $j$ is one of the lattice sites, and averaged over disorder realizations deep in the ETH regime of the $J_{1}$-$J_{2}$ model, is shown in Fig. 8(a). The spectral function possesses a characteristic structure of a plateau, which extends from the Heisenberg energy $\omega_{\mathrm{H}}$ at small frequencies to the Thouless energy $\omega_{\mathrm{Th}}$ at large $\omega$ [^22] Interested in scaling of $\omega_{Th}$ with system size $L$ and disorder strength $W$, we identify the notions of energy, angular speed and frequency.. These features motivated [582] to define the Thouless energy as $\omega$ at which the plateau develops. Here, the Thouless energy is extracted as a point at which $\ln\left(f^{2}_{0}/f^{2}(\omega))\right)>\epsilon$ where $\epsilon$ is a fixed threshold (we put $\epsilon=0.02$) and $f^{2}_{0}$ is the average of $f^{2}(\omega)$ over the plateau. Recalling that $f_{\mathcal{A}}(\bar{E},\omega)=1$ for the RMT, we note that appearance of the plateau marks the onset of RMT behavior of the spectral function $f^{2}(\omega)$. The large frequency tail of the spectral function $f^{2}(\omega)\propto\omega^{-1/2}$ corresponds to a diffusive decay of the autocorrelation function $\braket{S^{z}_{j}(t)S^{z}_{j}}$, which occurs deep in the ergodic regime of the $J_{1}$-$J_{2}$ model and lasts until the time scale proportional to $\omega_{\mathrm{Th}}^{-1}$. In principle, the Thouless time $t_{\mathrm{Th}}$ extracted from the behavior of SFF should scale in the same manner as $\omega_{\mathrm{Th}}^{-1}$ extracted here, reflecting the transport properties of the system. Approximately exponential scaling of $\omega_{\mathrm{Th}}\propto e^{-\kappa W}$ [572] with disorder strength is consistent with the behavior of the Thouless time extracted from the SFF [692].
@@
For a fixed system size $L$, the plateau of the spectral function $f^{2}(\omega)$ becomes narrower and the ratio $\omega_{\mathrm{Th}}/\omega_{\mathrm{H}}$ decreases with the growth of the disorder strength $W$. At the same time, the spectral function decays algebraically at high frequencies, $f^{2}(\omega)\propto\omega^{-\alpha}$, with a power $\alpha$ increasing with disorder strength $W$, reflecting the slow-down of system’s dynamics. At disorder strengths corresponding to the level statistics intermediate between the ETH and MBL regimes, the exponent $\alpha$ approaches values close to unity [416], which correspond either to a power-law decay of the autocorrelation function $\left\langle S^{z}_{j}(t)S^{z}_{j}\right\rangle\propto t^{-\beta}$ with a very small exponent $\beta$, or when $\alpha\to 1$, to a logarithmic decay time [572]. Upon further increase of the disorder strength, the averaged spectral function $f^{2}(\omega)$ starts to differ qualitatively from its typical counterpart. In the MBL regime, there are no traces of the plateau in $f^{2}(\omega)$, as shown in Fig. 8(b).
@@
Increase of the disorder strength leads to the accumulation of weight $C_{0}$ of the spectral function at zero frequency. In the limit $\omega\to 0$, the spectral function reproduces the infinite time average of the autocorelation function $\left\langle S^{z}_{j}(t)S^{z}_{j}\right\rangle$, reflecting the memory of the system about its initial state. The stiffness $C_{0}$ is shown in Fig. 8(b). In the ETH regime, $C_{0}$ decays exponentially with the system size $L$ indicating the ergodicity of system’s dynamics. In contrast, in the MBL regime, $C_{0}$ remains, to a good approximation, independent of $L$. Hence, $C_{0}$ follows the trends schematically represented in Fig. 3(b). The value of $C_{0}$ decreases with the growth of $L$, and this decrease is more pronounced with the increase of system size. This behavior of the results, similarly to the other examples discussed in this Section, does not allow for unambiguous conclusions about the fate of the system in the $L\to\infty$ limit.
@@
Understanding of evolution of the spectral function $f^{2}(\omega)$ with the disorder strength $W$, equivalent to comprehending the dynamics of autocorrelation functions $\left\langle A(t)A\right\rangle$ remain an outstanding challenge for the non-equilibrium physics of disordered many-body systems. A recent attempt [678] proposes to use the non-interacting limit ($\Delta=0$ in (7) or (27)) as the starting point. In the presence of interactions, Anderson LIOMs acquire finite relaxation times $t_{R}$. The distribution of $t_{R}$, relatable to the time dynamics of $\left\langle A(t)A\right\rangle$, becomes broad already for $W>W^{T}(L)$, i.e., directly above the ETH regime. At these disorder strengths, the off-diagonal elements of local observables show deviations from Gaussianity [387], while a fraction of Anderson LIOMs relaxation times $t_{R}$ exceeds the Heisenberg time $t_{H}$, see Sec. VII.3.1 for details. The latter operators appear to be exactly conserved at the considered system size, and the interplay between the broad distribution of $t_{R}$ and the exponentially increasing Heisenberg time determines the fate of the ETH-MBL crossover in the system.

### v3

@@ VI.2.2 Quantum information measures
$S_{A}=-\mathrm{tr}_{A}(\hat{\rho}_{A}\log_{2}\hat{\rho}_{A}),$ (38)
where $\hat{\rho}_{A}=\mathrm{tr}_{B}\ket{\psi}\bra{\psi}$ | is the reduced density matrix of subsystem $A$, obtained by tracing out the degrees of freedom of subsystem $B$. Quantum mutual information (QMI) [483] measures correlation between subsystems $A_{1}$ and $A_{2}$ of the system, and can be calculated as
@@
Entanglement entropy has been employed in the studies of MBL from its early days [753] and area-law for the entanglement entropy of eigenstates was proposed [55, 626]. It predicts that the entanglement entropy $S_{A}$ (38) is proportional to the boundary of the subsystem $A$, which, for 1D systems, implies $S_{A}\sim\mathrm{const}$ independently of the size $\ell_{A}$ of the subsystem $A$, provided that $\ell_{A}\gg 1$. This is consistent with the LIOMs picture, as discussed in Sec. III.4. In contrast, the mid-spectrum eigenstates in ergodic regime are nearly maximally entangled [496, 77], leading to volume-law $S_{A}\sim\ell_{A}$. Global conservation laws, such as $U(1)$ symmetry of (7), affect the constant term in the volume-law scaling of $S_{A}$ [735].
@@
For a meaningful comparison of entanglement entropy for different system sizes, it is convenient to rescale the entanglement entropy by the average $S_{\mathrm{RMT}}$ for Haar random states in the considered subspace of the Hilbert space, and consider $s=S_{A}/S_{\mathrm{RMT}}$. Averaging this quantity over mid-spectrum eigenstates |$\ket{m}$ and disorder realizations, leads to $\overline{s}$ which may serve as an ergodicity breaking indicator. The rescaled entanglement entropy, $\overline{s}$, across the ETH-MBL crossover in the $J_{1}$-$J_{2}$ model converges, with increasing the system size $L$, to unity in the ergodic regime and it decreases towards 00 in the MBL regime, as shown in Fig. 7(c). Even deep in the ergodic regime, we observe a small correction of the order of $\mathcal{O}(1)$ to the value of the averaged entanglement entropy [^30] For |$\ket{\psi}=U\ket{\psi_{0}}$ where |$\ket{\psi_{0}}$ is fixed and $U$ is drawn with Haar measure from the orthogonal group in the zero-magnetization sector, $\sum_{i}\hat{S}^{z}_{i}=0$, and the symmetric bipartition of the chain, one finds $S_{\mathrm{RMT}}=L/2-(1/(4\log(2))+1/2)$ [735]. The average entanglement entropy $S_{\mathrm{ERG}}$ of the mid-spectrum eigenstates of $J_{1}$-$J_{2}$ model in the ETH regime, e.g., at $W=1$, is well fitted by $S_{\mathrm{ERG}}=S_{\mathrm{RMT}}-c/\log(2)$, consistently with the arguments of [283], and $c\approx 0.10$, in approximate agreement with calculations of [302, 305]. . The magnitude of the correction depends on the choice of the parameters of the model [355]. Notably, we observe a significant drift of the crossing point of $\overline{s}(W)$ curves with the increase of system sizes. This drift is analogous to the behavior of the averaged gap ratio $\overline{r}$ and the fractal dimension $D_{q}$ discussed in the preceding sections, disallowing a single parameter scaling and preventing one from reaching unambiguous conclusions about the fate of the crossover in the $L\to\infty$ limit.
@@
Interestingly, the extent of the drift of the crossing point may depend on the considered observable. For instance, within the range of system sizes considered by us, the drifts of the crossing points are larger for the average gap ratio $\overline{r}$ and the rescaled entanglement entropy $\overline{s}$ than for the fractal dimension $D_{q}$, c.f. Fig. 4, Fig. 6, and Fig. 7, respectively. Similarly, the QMI calculated for extensive subsystems $A_{1}$ and $A_{2}$ of the Kicked Ising model in [645] exhibits a weaker drift of the crossing point than the entanglement entropy. A consistency of the extrapolations to the large $L$ limit of crossing points $W^{*}(L)$ yields consistent results for various observables [646, 643, 645]. The consistency between extrapolations of the crossing points for various observables may suggest their relevance. Nevertheless, such extrapolations remain uncontrolled in the absence of solid arguments for the form of the extrapolating functions.
@@ VI.3 Matrix elements of local observables
Matrix elements $\langle n|\hat{A}|m\rangle$ of local observables $\hat{A}$ in eigenstates |$\ket{m}$ of many-body Hamiltonian $\hat{H}$ provide a direct link between the features of system’s spectrum and measurable dynamical properties. The analysis of properties of matrix elements of local observables is an important way of quantifying ergodicity and its breakdown in many-body systems.
@@ VI.3.1 Spectral function
The ETH ansatz for the matrix elements $\langle n|\hat{A}|m\rangle$ (3) consists of the diagonal term ($m=n$) and the off-diagonal terms ($m\neq n$). We employ a spectral function of a local observable $\hat{A}$ as an organizing principle for studies of both diagonal and off-diagonal matrix elements of $\hat{A}$. The spectral function is defined as the Fourier transform of the autocorrelation function $\left\langle A(t)A\right\rangle$, i.e.,
@@
[…]
@@
[…]
@@
The $f^{2}(\omega)$ function is directly linked with the dynamics of the system as the Fourier transform of the autocorrelation function. The spectral function enters the fluctuation-dissipation relations [336], and, if the operator $\hat{A}$ is a density of a conserved quantity, e.g., a spin current operator [543], $f^{2}(\omega)$ is related to the conductivity of the system [8]. Comparison of (41) with the ETH ansatz provides a link between the spectral function and the structure function $f_{\mathcal{A}}(\bar{E},\omega_{mn})$ in the ETH ansatz (3). Moreover, the small frequency behavior of the spectral function, $f^{2}(\omega)\stackrel{{\scriptstyle\omega\ll 1}}{{=}}C_{0}\delta(\omega)$, reflects the averaged square of the diagonal matrix element of $\hat{A}$, $C_{0}=\frac{1}{|\mathcal{M}_{E_{c}}|}\sum_{m\in\mathcal{M}_{E_{c}}}|\bra{m}\hat{A}\ket{m}|^{2}$, since the length of the energy interval corresponding to $\mathcal{M}_{E_{c}}$ vanishes in the large $L$ limit.
@@
Figure 8: The spectral function $f^{2}(\omega)$ calculated for $\hat{A}=\hat{S}^{z}_{j}$ in the $J_{1}$-$J_{2}$ spin chain of size $L$ in the ergodic regime, (a), for $W=1.5$ and in the MBL regime, (b), for $W=12$. The Heisenberg energy $\omega_{\mathrm{H}}\sim e^{-cL}$ is denoted by red squares, while the Thouless energy $\omega_{\mathrm{Th}}$ by red circles. The dashed line in (a) denotes a diffusive tail $f^{2}(\omega)\propto\omega^{-1/2}$. (c) The stiffness $C_{0}$ as function of disorder strength $W$ for various $L$ distinguishes ETH and MBL regimes.
@@
The spectral function $f^{2}(\omega)$ evaluated for the operator $\hat{A}=\hat{S}^{z}_{j}$, where $j$ is one of the lattice sites, and averaged over disorder realizations deep in the ETH regime of the $J_{1}$-$J_{2}$ model, is shown in Fig. 8(a). The spectral function possesses a characteristic structure of a plateau, which extends from the Heisenberg energy $\omega_{\mathrm{H}}$ at small frequencies to the Thouless energy $\omega_{\mathrm{Th}}$ at large $\omega$ [^32] Interested in scaling of $\omega_{Th}$ with system size $L$ and disorder strength $W$, we identify the notions of energy, angular speed and frequency.. These features motivated [629] to define the Thouless energy as the value of $\omega$ at which the plateau develops. Here, the Thouless energy is extracted as a point at which $\ln\left(f^{2}_{0}/f^{2}(\omega))\right)>\epsilon$ where $\epsilon$ is a fixed threshold (we put $\epsilon=0.02$) and $f^{2}_{0}$ is the average of $f^{2}(\omega)$ over the plateau. Recalling that $f_{\mathcal{A}}(\bar{E},\omega)=1$ for the RMT, we note that the appearance of the plateau marks the onset of RMT behavior of the spectral function $f^{2}(\omega)$. The large frequency tail of the spectral function, $f^{2}(\omega)\propto\omega^{-1/2}$, corresponds to a diffusive decay of the autocorrelation function $\braket{\hat{S}^{z}_{j}(t)\hat{S}^{z}_{j}}$, which occurs deep in the ergodic regime of the $J_{1}$-$J_{2}$ model and lasts until the time scale proportional to $\omega_{\mathrm{Th}}^{-1}$. In principle, the Thouless time $t_{\mathrm{Th}}$ extracted from the behavior of SFF should scale in the same manner as $\omega_{\mathrm{Th}}^{-1}$ extracted here, reflecting the transport properties of the system. Approximately exponential scaling of $\omega_{\mathrm{Th}}$ with the disorder strength, $\omega_{\mathrm{Th}}\propto e^{-\kappa W}$ [619], is consistent with the behavior of the Thouless time extracted from the SFF [748].
@@
For a fixed system size $L$, the plateau of the spectral function $f^{2}(\omega)$ becomes narrower and the ratio $\omega_{\mathrm{Th}}/\omega_{\mathrm{H}}$ decreases with the growth of the disorder strength $W$. At the same time, the spectral function decays algebraically at high frequencies, $f^{2}(\omega)\propto\omega^{-\alpha}$, with a power $\alpha$ increasing with disorder strength $W$, reflecting the slow-down of system’s dynamics. At disorder strengths corresponding to the level statistics intermediate between the ETH and MBL regimes, the exponent $\alpha$ approaches values close to unity [452], which correspond either to a power-law decay of the autocorrelation function $\left\langle\hat{S}^{z}_{j}(t)\hat{S}^{z}_{j}\right\rangle\propto t^{-\beta}$, with a very small exponent $\beta$, or when $\alpha\to 1$, to a logarithmic decay time [619]. Upon further increase of the disorder strength, the averaged spectral function $f^{2}(\omega)$ starts to differ qualitatively from its typical counterpart. In the MBL regime, there are no traces of the plateau in $f^{2}(\omega)$, as shown in Fig. 8(b).
@@
Increase of the disorder strength leads to the accumulation of weight $C_{0}$ of the spectral function at zero frequency. In the limit $\omega\to 0$, the spectral function reproduces the infinite time average of the autocorelation function $\left\langle\hat{S}^{z}_{j}(t)\hat{S}^{z}_{j}\right\rangle$, reflecting the memory of the system about its initial state. The stiffness $C_{0}$ is shown in Fig. 8(b). In the ETH regime, $C_{0}$ decays exponentially with the system size $L$ indicating the ergodicity of system’s dynamics. In contrast, in the MBL regime, $C_{0}$ remains, to a good approximation, independent of $L$. Hence, $C_{0}$ follows the trends schematically represented in Fig. 3(b). This behavior follows the general trend illustrated in Fig. 3 (b) and, similarly to the other examples discussed in this Section, does not allow for unambiguous conclusions about the fate of the system in the $L\to\infty$ limit.
@@
Understanding of evolution of the spectral function $f^{2}(\omega)$ with the disorder strength $W$, equivalent to comprehending the dynamics of autocorrelation functions $\left\langle\hat{A}(t)\hat{A}\right\rangle$, remains an outstanding challenge for the non-equilibrium physics of disordered many-body systems. A recent attempt [733] proposes to use the non-interacting limit ($\Delta=0$ in (7) or (27)) as the starting point. In the presence of interactions, Anderson LIOMs acquire finite relaxation times $t_{R}$. The distribution of $t_{R}$, relatable to the dynamics of $\left\langle\hat{A}(t)\hat{A}\right\rangle$, becomes broad already for $W>W^{T}(L)$, i.e., directly above the ETH regime. At these disorder strengths, the off-diagonal elements of local observables show deviations from Gaussianity [423], while a fraction of Anderson LIOMs relaxation times $t_{R}$ exceeds the Heisenberg time $t_{H}$, see Sec. VII.3.1 for details. The latter operators appear to be exactly conserved at the considered system size, and the interplay between the broad distribution of $t_{R}$ and the exponentially increasing Heisenberg time determines the fate of the ETH-MBL crossover in the system.